Answer graphically
2x+3y = 0
x+5y = 7

Answer by mananth(16949)

2x+3y = 0.......................(1)
x+5y = 7........................(2)

Find the Y & X intercepts of the two lines by plugging x=0, & y=0
equation1
when x= 0, y= 0 ( 0, , 0 )
when y = 0, x= 0 ( 0 , 0, )
Equation 2
when x= 0, y= 1.4 ( 0, , 1.4 )
when y = 0, x= 7 ( 7 , 0, )

The solution is (-3,2)